OpenStudy (alex_mattucci):

Help! I need to factor 2x^2 + bx + 30 but I need a number for b!

OpenStudy (alex_mattucci):

@SolomonZelman

OpenStudy (anonymous):

let b=a+c such that a*c=60 b=10+6=16 10*6=60 60=2*30 =3*20 =4*15 =5*12 =6*10 only possibility is 10 and 6 or -10-6 -10*-6=60 -10-6=-16 so b is 16 or-16

OpenStudy (anonymous):

number is only 16 and integers are -16 and 16

OpenStudy (alex_mattucci):

ok, thank you so much!

OpenStudy (anonymous):

yw